12. For a school band trip, Marcia decides to order T-shirts for all of the participants. It will cost $4 per shirt for the medium size, and $5 per shirt for the large size. Marcia orders a total of 70 T-shirts and spends $320. How many are medium shirts?
(the answer is: 30 medium T-shirts, but i dont know how to solve it?)
no i dont know how to do it via.

Answer by ankor@dixie-net.com(22740)

12. For a school band trip, Marcia decides to order T-shirts for all of the
participants. It will cost $4 per shirt for the medium size, and $5 per shirt
for the large size. Marcia orders a total of 70 T-shirts and spends $320.
How many are medium shirts?
:
Let L = no. of lg shirts
Let M = no. of med shirts:
:
Write an equation for the total number of shirts
L + M = 70
or
L = (70-M); we can use this for substitution
:
Write an equation for the total cost of the shirts
5L + 4M = 320
:
From the total shirts equation, substitute (70-M) for L in the above equation
5(70-M) + 4M = 320
:
350 - 5M + 4M = 320
:
-1M = 320 - 350
-M = -30
which is
M = 30 shirts
;
:
Do you think you can do this kind of problem now? Any questions?